4mg of NaCl was weighed and topped up to 50ml.

Calculate the concentration of Na in ppm


1
Expert's answer
2021-06-28T08:11:25-0400

"ppm = \frac{mass \;of\; Na^+}{mass\; of \;solvent} \times 10^6"

Proportion:

58.5 g NaCl contain 23 g Na+

"4 \times 10^{-3}" g NaCl contain x g Na+

"x = \frac{4 \times 10^{-3} \times 23}{58.5}=1.57 \times 10^{-3} \;g \\

ppm = \frac{1.57 \times 10^{-3}}{50} \times 10^6 = 31.45 \;ppm"

Answer: 31.45 ppm